# Interpreting CLIP with Sparse Linear

### v1

@@ Abstract
footnotetext: $*{}^{*}$ Equal contribution, order by coin flip.footnotetext: ${}^{\dagger}$ Equal contribution, alphabetic order.footnotetext: ${}^{a}$ Kempner Institute for the Study of Natural & Artificial Intelligencefootnotetext: ${}^{b}$ School of Engineering and Applied Sciencesfootnotetext: ${}^{c}$ Harvard Business School
@@ 1 Introduction
Natural images contain complex semantic information, such as the objects they contain, the scenes they depict, the actions performed, and the relationships between them. Machine learning models on visual data aim to encode this semantic information in their representations, to help with a wide variety of downstream tasks, such as object classification, scene recognition, or action recognition. However, it is often difficult to enforce explicit encoding of these semantics within model representations, and even harder to interpret these semantics post hoc. Further, these model representations can be brittle, encoding idiosyncratic patterns specific to individual datasets, as opposed to high-level semantic information. Multimodal models have been proposed as a potential solution to this brittleness issue, and methods such as CLIP Radford et al. (2021) have empirically been found to provide highly performant, semantically rich representations of image data. The richness of these representations is evident from their high performance on a wide variety of tasks, such as zero-shot classification and image retrieval Radford et al. (2021), image captioning Mokady et al. (2021), and image generation Ramesh et al. (2022).However, despite their performance, it remains unclear how to quantify the semantic content contained in their dense representations. In this work, we answer the question: how can we interpret CLIP embeddings in terms of the semantics of the underlying data they encode?
Existing machine learning literature in areas such as concept bottleneck models Koh and Liang (2017), disentangled representation learning Bengio (2013), and mechanistic intepretability Olah et al. (2017) have proposed various approaches to understanding the semantics behind representations. However, these methods generally require a predefined set of concepts Kim et al. (2018), data with concept labels Hsu et al. (2023), or rely on qualitative visualizations, which can be unreliable Geirhos et al. (2023). Similar to these lines of work, we aim to recover representations that reflect the underlying semantics of the inputs. However, distinct from these works, we propose to do this in a post hoc fashion, without concept labels, training, or qualitative analysis of visualizations.
In this work, we leverage the highly structured and multimodal nature of CLIP embeddings for interpretability, and propose SpLiCE , which decomposes CLIP representations into a set of human-interpretable, semantically meaningful concepts. Our overall contributions are:
  • In Sections 3 and 4, we formalize the sufficient conditions under which decomposition is feasible, and then introduce SpLiCE , a novel method that decomposes dense CLIP embeddings into sparse combinations of human-interpretable, semantic concepts.
  • Our extensive experiments in Section 5 reveal that SpLiCE recovers highly sparse111we recommend and use sparsity levels of $\sim$ 10-30 in practice, interpretable representations with minimal loss in performance on downstream tasks, while accurately capturing the semantics of the underlying inputs.
  • In Section 6, we present three case studies for applying SpLiCE : spurious correlation detection, model editing, and quantifying distribution shifts in datasets. We find that SpLiCE is able to uncover unknown biases in datasets, such as the problematic prevalence of women in bikinis in CIFAR100 and the out-of-distribution increase in yellow cars in the early 2000s in the Stanford Cars dataset, while also allowing for intervention on such correlations via representation and model editing.
@@ 2 Related Work
Linear Representation Hypothesis. In language modeling, the linear representation hypothesis suggests that semantic concepts are organized linearly in the latent space of a model Mikolov et al. (2013); Park et al. (2023); Arora et al. (2018, 2016); Faruqui et al. (2015). Recent work has also shown that multimodal models behave like bags-of-words, independent of input sequence order Yuksekgonul et al. (2022a). Merullo et al. (2022) and Seth et al. (2023) show that there exists a linear mapping between image and text embeddings in arbitrary models. Motivated by these findings, we show that the linear representation hypothesis exists in a modality-agnostic semantic space, allowing for a translation between dense, modality-specific embeddings and sparse, semantic representations.
@@
Figure 2: Example images from MSCOCO shown with their captions below and their concept decompositions on the right. We display the top six concepts for visualization purposes, but images in the figure had decompositions with 7-20 concepts.
Concept-based Explanations. Concept Bottleneck Models (CBMs) Koh et al. (2020), and earlier work on attribute based models Lampert et al. (2009); Torresani et al. (2010); Kumar et al. (2009) learn an intermediate representation of scores over concepts for use with a final linear layer, creating interpretable representations that can be intervened on to correct biases and mistakes in models. However, CBMs require training constrained architectures with concept-labeled data, and recent work has shown that CBMs can learn to leak downstream task information through the concept bottleneck layer and may fail to satisfy notions of interpretability and utility Margeloiu et al. (2021); Mahinpei et al. (2021).
Mechanistic Interpretability and Disentanglement. Mechanistic interpretability explains representations through model activations, by labelling circuits and neurons in networks via feature visualization Olah et al. (2017, 2020), or by measuring concept activations and directions in latent space Bau et al. (2017); Fong and Vedaldi (2018); Kim et al. (2018); McGrath et al. (2022); Lucieri et al. (2020); Zhou et al. (2018). Fel et al. (2023) combine these methods, using dictionary learning to extract visual concept activations, whose semantics can be identified via feature visualization.
Work in disentangled representation learning has developed architectures that capture independent factors of variation in data Hsu et al. (2023); Bricken et al. (2023); Chen et al. (2018); Bengio (2013); Comon (1994); Hyvärinen and Oja (2000), allowing for manual probing of disentangled representations for human-interpretable concepts. While promising, these methods all rely on labelled concept sets, manual labelling of visualizations, or computationally intensive exhaustive searches over networks. Furthermore, Geirhos et al. (2023); Makelov et al. (2023) showcase the unreliability of feature visualization and activation patching, respectively.
CLIP Interpretability. The works most methodologically similar to ours leverage the semantic structure of CLIP and its text encoder to provide explanations. Moayeri et al. (2023), Yuksekgonul et al. (2022b), and Yun et al. (2022) construct concept similarity scores of image embeddings for use by downstream CBMs or probes, but these scores do not provide semantic decompositions of the original embeddings, and are not interpretable due to their lack of sparsity and nonnegativity. Chen et al. (2023) train a custom vision-language architecture with a sparse latent dictionary, but their method relies on properties of transformer backbones and cannot be used post-hoc to explain existing models. Gandelsman et al. (2023) also leverage the text encoder of CLIP to explain components of the image embedding, but are limited to ViT architectures and take a mechanistic interpretability-style approach requiring a probe dataset of text. Finally, Chattopadhyay et al. (2023) use orthogonal matching pursuit to solve the a similar sparse dictionary learning problem over CLIP image embeddings and text explanation embeddings, but do not impose non-negativity and use complex explanations that do not capture the individual semantic units making up an image.
@@ 3 When do Sparse Decompositions Exist?
In this section, we aim to answer the question: under what conditions can CLIP representations be decomposed into sparse semantic representations? To address this question, we are required to reason about both the properties of CLIP as well as the properties of the underlying data.
@@
Notation. Let $\mathbf{x}^{\mathrm{img}}\in\mathbb{R}^{d_{i}}$, $\mathbf{x}^{\mathrm{txt}}\in\mathbb{R}^{d_{t}}$ be image and text data respectively. Given CLIP image encoder $f:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d}$ and text encoder $g:\mathbb{R}^{d_{t}}\rightarrow\mathbb{R}^{d}$, we define CLIP representations in $\mathbb{R}^{d}$ as $\mathbf{z}^{\mathrm{img}}=f(\mathbf{x}^{\mathrm{img}})$ and $g(\mathbf{x}^{\mathrm{txt}})=\mathbf{z}^{\mathrm{txt}}$. Our method uses dictionary learning to approximate $\mathbf{z}^{\mathrm{img}}$ with a concept decomposition $\mathbf{w}^{*}\in\mathbb{R}_{+}^{c}$ over a fixed concept vocabulary $\mathbf{C}\in\mathbb{R}^{d\times c}$. We define the resulting reconstruction of $\mathbf{z}^{\mathrm{img}}$ from $\mathbf{C}$ and $\mathbf{w}^{*}$ as $\hat{\mathbf{z}}^{\mathrm{img}}$.
@@
The overall goal of our method is to approximate $f(\mathbf{x}^{\text{img}})\approx\mathbf{C}\mathbf{w}^{*}$, such that $\mathbf{w}^{*}$ is non-negative and sparse, and in this section we formalize when this is possible. We begin by considering a data-generating process for coupled image and text samples. Specifically, we model the generative process parameterized by a $k$-dimensional latent concept vector ${\omega}\in\mathbb{R}_{+}^{k_{1}}$ and a random vector $\epsilon\in\mathbb{R}^{k_{2}}$ as
$\mathbf{x}^{\text{img}}={h}^{\text{img}}(\omega,\epsilon),~{}~{}~{}~{}\mathbf{x}^{\text{txt}}=h^{\text{txt}}(\omega,\epsilon),$
$\omega\sim\rho,$
$\epsilon\sim\phi,$
@@
where $\rho$ is a prior distribution over semantic concepts, $\phi$ is a prior distribution over nonsemantic concepts (such as camera orientation and lighting for images or sentence structure for text), and $h^{\text{img}}:\mathbb{R}^{k_{1}+k_{2}}\rightarrow\mathbb{R}^{d_{i}}$, and $h^{\text{txt}}:\mathbb{R}^{k_{1}+k_{2}}\rightarrow\mathbb{R}^{d_{t}}$ are mappings from latent variables $(\omega,\epsilon)$ to data. In addition, we assume unique invertibility, i.e., for every $\mathbf{x}^{\text{img}},\mathbf{x}^{\text{txt}}$, there exists a unique pre-images $(\omega,\epsilon)$ that generated it. Here, each coordinate $\omega_{i}\in\mathbb{R}_{+}$ encodes the degree of prevalence of the $i^{\text{th}}$ concept in the underlying data, where a concept is a semantic piece of information that can be described in a single English word. Under this generative model, a total of $k_{1}$ distinct concepts can be encoded by inputs, one corresponding to each coordinate of $\omega$.
Note that these data-generating processes and latent concept vectors cannot be directly observed. We now identify a key property, the sparsity of data in concept space.
Property 3.1.
@@
Data is sparse in concept space. In other words, for some $\alpha\ll k_{1}$, we have $\|\omega\|_{0}\leq\alpha,\forall~{}\omega\sim\rho$.
The next property is the most critical one and posits that CLIP models are linear over concepts, which is also closely related to the linear representation hypothesis Park et al. (2023). In addition, we posit that CLIP is largely invariant to the ‘non-semantic’ factors $\epsilon$ in the image. For example, the embedding of a cat photographed from two angles should be the same. To formalize this in our context, recall that the image and text encoders of CLIP are given by $f$ and $g$, respectively.
Now consider $f(\mathbf{x}^{\text{img}})=f(h^{\text{img}}(\omega,\epsilon))\triangleq(f\circ h^{\text{img}})(\omega,\epsilon)$, and similarly $g(\mathbf{x}^{\text{txt}})=(g\circ h^{\text{txt}})(\omega,\epsilon)$. Note that while these composition maps $f\circ h^{\text{img}}$ and $g\circ h^{\text{txt}}$ can be nonlinear in general, the following property indicates that these are in fact linear functions of $\omega$, and independent of $\epsilon$.
Property 3.2.
CLIP encoders are (1) independent of $\epsilon$, i.e.,
@@
$g\circ h^{\text{img}}(\omega)$ $\triangleq g\circ h^{\text{img}}(\omega,\epsilon),$
$g\circ h^{\text{txt}}(\omega)$ $\triangleq g\circ h^{\text{txt}}(\omega,\epsilon),$
and (2) linear in concept space $\omega$, i.e., $g\circ h^{\text{txt}}$ and $f\circ h^{\text{img}}$ are linear wrt $\omega$.
Note that this assumption does not require $f,g$ to be linear, rather it requires its composition with the underlying generative model $h^{\text{img}},h^{\text{txt}}$ to result in a linear map. This implies that the CLIP encoders $f,g$ perform an “inversion” of the latent generative process, up to a linear transformation. This assumption falls in line with the idea that computer vision systems must behave like “inverse graphics” models Hinton (2013); Romaszko et al. (2017). In Appendix E.1, we provide evidence for such linearity of CLIP in a toy setting.
We now make another assumption connecting the image and text representations of CLIP, requiring them to be equal. This is approximately enforced by CLIP’s training objective, which requires image and text representations to align.
Property 3.3.
For the same latent concept vector $\omega$, the CLIP image and text representations are equal to each other, i.e., $f\circ h^{\text{img}}(\omega)=g\circ h^{\text{txt}}(\omega)$.
Our final assumption involves the computation of the basis vectors in concept space. These are given by $\mathbf{e}_{i}\in\mathbb{R}^{k}$, which are unique one-hot vectors with a one at the $i^{\text{th}}$ coordinate. Consider the mapping of these “basis” concepts in text space. As concepts can are described by a single English word without any grammatical structure or phrasing involved, the mapping of a basis vector to text is independent of non-semantic noise. The following assumption states that these are text strings of individual words depicting that concept.
Property 3.4.
Basis concepts $\mathbf{e}_{i}$ map to individual words in text space, i.e., $h^{\text{txt}}(\mathbf{e}_{i},\epsilon)=h^{\text{txt}}(\mathbf{e}_{i})$ is a text string of an individual word representing that concept.
[…]
Theorem 3.5.
Given Properties 4.1 - 4.4, CLIP image embeddings $f$ can be written as a sparse linear combination of text embeddings, i.e,
$f(\mathbf{x}^{\mathrm{img}})=\mathbf{C}\mathbf{w};~{}~{}s.t.~{}~{}\|\mathbf{w}\|_{0}\leq\alpha$
where $\mathbf{w}\in\mathbb{R}_{+}^{k_{1}}$, and $\mathbf{C}\in\mathbb{R}^{d\times k_{1}}$, which is the concept dictionary obtained from word embeddings of individual words, as described above.
Proof Sketch.
Property 3.1 + 3.2 together show that $f$ can be written as a sparse linear combination of image embeddings, and Property 3.3 + 3.4 help transfer this result to the text embeddings of individual words. We provide the full proof in Appendix C. ∎
Thus we have shown that given natural assumptions about CLIP and the data generating process in terms of Properties 3.1–3.4, CLIP models have a sparse decomposition property. The critical idea here is Property 3.2, which requires models to behave linearly in concept space. In the next section, we propose a method to perform this sparse decomposition.

### v2

@@ Abstract
footnotetext: Equal contribution, order by coin flip.    Equal contribution, alphabetic order.footnotetext: a Kempner Institute for the Study of Natural & Artificial Intelligencefootnotetext: b School of Engineering and Applied Sciences   c Harvard Business School
@@ 1 Introduction
Natural images include complex semantic information, such as the objects they contain, the scenes they depict, the actions being performed, and the relationships between them. Machine learning models trained on visual data aim to encode this semantic information in their representations to perform a wide variety of downstream tasks, such as object classification, scene recognition, segmentation, or action prediction. However, it is often difficult to enforce explicit encoding of these semantics within model representations, and it is even harder to interpret these semantics post hoc to better understand what models may have learnt and how they leverage this information. Further, model representations can be brittle, encoding idiosyncratic patterns specific to individual datasets and modalities instead of general human-interpretable semantic information. Multimodal models have been proposed as a potential solution to this issue, and methods such as CLIP [1] have empirically been found to provide highly performant, semantically rich representations of image data. The richness of these representations is evident from their high performance on a variety of tasks, such as zero-shot classification and image retrieval [1], image captioning [2], and image generation [3]. However, despite their performance, it remains unclear how to quantify the semantic content contained in their dense representations. In this work, we answer the question: can we decompose CLIP embeddings into human-interpretable representations of the semantic concepts they encode? This can provide insight into the types of tasks CLIP can solve, the biases it may contain, and the manner through which downstream predictions are made.
Existing literature in areas such as concept bottleneck models [4], disentangled representation learning [5], and mechanistic intepretability [6] have proposed various approaches to understanding the semantics encoded by representations. However, these methods generally require predefined sets of concepts [7], data with concept labels [8], or rely on qualitative visualizations, which can be unreliable [9]. Similar to these lines of work, we aim to recover representations that reflect the underlying semantics of the inputs. However, distinct from these works, we propose to do this in a task-agnostic manner and without concept datasets, training, or qualitative analysis of visualizations.
Our method, SpLiCE , leverages the highly structured and multimodal nature of CLIP embeddings for interpretability, and decomposes CLIP representations via a semantic basis to yield a sparse, human-interpretable representation. Remarkably, these interpretable SpLiCE embeddings have favorable accuracy-interpretability tradeoffs when compared to black-box CLIP representations on metrics such as zero-shot accuracy. Our overall contributions are:
@@
  • In Sections 3 and 4, we formalize the sufficient conditions under sparse decomposition of CLIP is feasible, and introduce SpLiCE , a novel method that decomposes dense CLIP embeddings into sparse, human-interpretable concept embeddings.
  • Our extensive experiments in Section 5 reveal that SpLiCE recovers highly sparse111we recommend and use sparsity levels of $\sim$ 10-30 in practice, interpretable representations with high performance on downstream tasks, while accurately capturing the semantics of the underlying inputs.
  • In Section 6, we present two case studies for applying SpLiCE : spurious correlation detection, and model editing. Using SpLiCE , we uncover a spurious correlation in the CIFAR100 dataset, where we find the "woman" concept and the "swimwear" concept to be correlated owing to the prevalence of women in swimwear in CIFAR100.
@@ 2 Related Work
Linear Representation Hypothesis. In language modeling, the linear representation hypothesis suggests that many semantic concepts are approximately linear functions of model representations [10, 11, 12, 13, 14]. Recent work has also shown that multimodal models encode concepts additively, behaving like bags-of-words representations [15]. Relatedly, [16, 17] show that there exists a linear mapping between image and text embeddings in arbitrary models. Our work makes use of these distinct but related observations to convert dense CLIP representations to sparse semantic ones.
@@
Figure 2: Example images from MSCOCO shown with their captions below and their concept decompositions on the right. We display the top seven concepts for visualization purposes, but images in the figure had decompositions with 7-20 concepts.
Concept Bottlenecks and Attribute Learning. Concept Bottleneck Models (CBMs) [18], and attribute-based models [19, 20, 21] learn intermediate representations of scores over concepts or image attributes for use with a final linear classification head, creating interpretable concept representations. However, these require expert-labeled concept or attribute datasets to train, which is expensive. Recent work on concept-bottlenecks for multimodal models avoids needing such labeled datasets, but still requires concept labels for specific tasks, which is obtained by querying large language models (LLMs) [22, 23, 24], making these methods task-specific and heavily reliant on the domain knowledge and subject to the biases of LLMs. On the other hand, SpLiCE uses a large-scale and overcomplete concept dictionary, avoiding dependence on training, specific domain knowledge, or a downstream task. Consequently, it can even be applied to understand unstructured, unsupervised image datasets in a label-free manner.
Mechanistic Interpretability and Disentanglement. Mechanistic interpretability explains representations through model activations, by labeling circuits and neurons in networks with feature visualization [6, 25] or by measuring concept activations and directions in latent space [26, 27, 7, 28, 29, 30]. Recent work [31] combines these methods, using dictionary learning to extract visual concept activations, whose semantics can be identified via feature visualization. Work in disentangled representation learning has developed architectures that capture independent factors of variation in data [8, 32, 33, 5, 34, 35], allowing for manual probing of disentangled representations for human-interpretable concepts. In both mechanistic interpretability and disentangled representation learning, methods typically rely on labeled concept sets, manual labeling of visualizations, or computationally intensive searches over data and latent representations or neurons to identify concepts. However, associating human-interpretable semantics with arbitrary neurons or latent directions is challenging, leading to the unreliability [9, 36] exhibited by such methods. Our approach side-steps this issue by decomposing CLIP representations into a predetermined set of concepts.
CLIP Interpretability. Many recent works leverage the semantic structure of CLIP and its text encoder to interpret representations. For example, [37], [38], and [39] construct concept similarity scores of image embeddings for use by downstream CBMs or probes, but these representations are not interpretable due to their lack of sparsity and the presence of negative concepts. chen2023stair create a custom vision-language architecture with a sparse latent dictionary, but it requires training from scratch and cannot be used post-hoc to explain existing models. Gandelsman et al. [40] also leverage the text encoder of CLIP to explain components of the image embedding, but are limited to ViT architectures and take a mechanistic interpretability-style approach requiring a labeled text dataset. Chattopadhyay et al. [22] build concept bottlenecks for specific classification tasks by expressing CLIP image representations as a sparse linear combination of task-specific concept vectors. However, their decomposition includes negative concepts, reducing interpretability, and uses task-specific concept dictionaries. Grootendorst [41] generate textual topics of datasets through multimodal topic modeling, which cannot provide explanations of individual representations. Distinct from these works, SpLiCE is more interpretable due to its sparsity, overcompleteness, and non-negativity, and is task-agnostic, aiming to serve as a drop-in replacement for black-box CLIP representations without requiring training.
@@ 3 When do Sparse Decompositions Exist?
In this section, we aim to answer the question: under what conditions can CLIP representations be decomposed into sparse semantic representations? To do so, we must reason about both the properties of CLIP as well as the properties of the underlying data.
@@
Notation. Let $\mathbf{x}^{\mathrm{img}}\in\mathbb{R}^{d_{i}}$, $\mathbf{x}^{\mathrm{txt}}\in\mathbb{R}^{d_{t}}$ be image and text data, respectively. Given the CLIP image encoder $f:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d}$ and text encoder $g:\mathbb{R}^{d_{t}}\rightarrow\mathbb{R}^{d}$, we define CLIP representations in $\mathbb{R}^{d}$ as $\mathbf{z}^{\mathrm{img}}=f(\mathbf{x}^{\mathrm{img}})$ and $\mathbf{z}^{\mathrm{txt}}=g(\mathbf{x}^{\mathrm{txt}})$. Our method uses dictionary learning to approximate $\mathbf{z}^{\mathrm{img}}$ with a concept decomposition $\mathbf{w}^{*}\in\mathbb{R}_{+}^{c}$ over a fixed concept vocabulary $\mathbf{C}\in\mathbb{R}^{d\times c}$. We define the resulting reconstruction of $\mathbf{z}^{\mathrm{img}}$ from $\mathbf{C}$ and $\mathbf{w}^{*}$ as $\hat{\mathbf{z}}^{\mathrm{img}}$.
@@
The goal of our method is to approximate $f(\mathbf{x}^{\text{img}})\approx\mathbf{C}\mathbf{w}^{*}$, such that $\mathbf{w}^{*}$ is non-negative and sparse, and in this section we formalize when this is possible. We begin by considering a data-generating process for coupled image and text samples. Specifically, we model the generative process parameterized by a $k$-dimensional latent concept vector ${\omega}\in\mathbb{R}_{+}^{k}$ and a random noise vector $\epsilon\in\mathbb{R}^{l}$ as
$\mathbf{x}^{\text{img}}={h}^{\text{img}}(\omega,\epsilon),~{}~{}~{}~{}\mathbf{x}^{\text{txt}}=h^{\text{txt}}(\omega,\epsilon),~{}~{}~{}\omega\sim\rho,~{}~{}~{}\epsilon\sim\phi,$
@@
where $\rho$ is a prior distribution over semantic concepts, $\phi$ is a prior distribution over nonsemantic concepts (such as camera orientation and lighting for images or arbitrary choices between synonyms for text), and $h^{\text{img}}:\mathbb{R}^{k+l}\rightarrow\mathbb{R}^{d_{i}}$, and $h^{\text{txt}}:\mathbb{R}^{k+l}\rightarrow\mathbb{R}^{d_{t}}$ represent the real-world data-generating process from latent variables $(\omega,\epsilon)$ to images and text respectively. Here, each coordinate $\omega_{i}\in\mathbb{R}_{+}$ encodes the degree of prevalence of the $i^{\text{th}}$ concept in the underlying data. We now list a set of sufficient conditions for our data-generating process and CLIP that admit a sparse decomposition of images into concepts.
Sufficient Conditions for Sparse Decomposition.
  1. 1.
@@
    Images and text are sparse in concept space: for some $\alpha\ll k$, we have $\|\omega\|_{0}\leq\alpha,\forall~{}\omega\sim\rho$.
  2. 2.
    CLIP captures semantic concepts $\omega$ and not $\epsilon$: $\forall\epsilon,\epsilon^{\prime},f\circ h^{\text{img}}(\omega,\epsilon)=f\circ h^{\text{img}}(\omega,\epsilon^{\prime})$ and similarly for $h^{\text{txt}}.$
  3. 3.
    CLIP is linear in concept space: $g\circ h^{\text{txt}}$ and $f\circ h^{\text{img}}$ are linear in $\omega$.
  4. 4.
@@
    CLIP image and text encoders are aligned: for a given $\omega$, $f\circ h^{\text{img}}(\omega,\epsilon)=g\circ h^{\text{txt}}(\omega,\epsilon)$.
We emphasize that the goal of enumerating a set of sufficient conditions for sparse decomposition is not to claim that these exactly hold in practice, but rather to reason about when sparse decompositions–as done in this work–are appropriate. In the Appendix (Section A.1, Prop. 1) we formalize and prove this claim, but in the interest of simplicity we keep the discussion here informal. We note that many of these are natural; Assumption 1 reflects how real-world images and text are simple and rarely contain complex semantic content, and the CLIP training process optimizes for Assumption 2 and 4222In practice we find that CLIP’s image and text encoders are not fully aligned, so we apply a preprocessing step (Sec 4.1).. Of these, the most critical one is Assumption 3, which closely relates to the linear representation hypothesis [11], which we investigate below.
Sanity Checking CLIP’s Linearity.
Table 1: Sanity checking the linearity of CLIP Embeddings.
$w_{a}$ $w_{b}$ cosine($\hat{z},z$)
ImageNet 0.48 $\pm$ 0.09 0.45 $\pm$ 0.09 0.76 $\pm$ 0.05
CIFAR100 0.45 $\pm$ 0.08 0.42 $\pm$ 0.08 0.75 $\pm$ 0.03
MIT States 0.48 $\pm$ 0.09 0.45 $\pm$ 0.09 0.76 $\pm$ 0.05
COCO Text 0.59 $\pm$ 0.12 0.47 $\pm$ 0.12 0.88 $\pm$ 0.04
We provide evidence for the third assumption, the linearity of CLIP, in a toy setting. We begin by asking the following question to confirm the general linearity of CLIP embeddings: “if two inputs are concatenated, does their joint embedding equal the average of their two individual embeddings?". For the image domain, we combine two images, $x_{a},x_{b}$, to form their composition $x_{ab}$ by placing $x_{a}$ in the top left quarter and $x_{b}$ in the bottom right quarter of a blank image. For the text domain, we simply append text $x_{b}$ to text $x_{a}$ to form $x_{ab}$. We then embed $x_{a},x_{b},x_{ab}$ with CLIP to get $z_{a},z_{b},z_{ab}$. Solving the equation $w_{a}*z_{a}+w_{b}*z_{b}=z_{ab}$ for scalar weights $w_{a},w_{b}$ then allows us to assess the linearity of $z_{a},z_{b},z_{ab}$. We report $w_{a},w_{b}$ and the cosine similarity between $\hat{z}_{ab}=[z_{a},z_{b}]\cdot[w_{a},w_{b}]$ and $z_{ab}$ in Table 1.
In general, we find that the composition of two inputs results in an embedding that is approximately equal to the average of the two input components, with $w_{a},w_{b}$ being very close to 0.5 across all datasets and for both modalities, providing preliminary evidence for the linearity of CLIP embeddings for both image and language.